OpenStudy (anonymous):

The equation of line LM is y = 5x + 4. Write an equation of a line perpendicular to line LM in slope-intercept form that contains point (-3, 2).

OpenStudy (anonymous):

Select one: a. y = -5x + 13 b. y= -1/5x + 7/5 c. y= 1/5x - 7/5 d. y= 5x-17

OpenStudy (anonymous):

nurali can you help me please on some questions

OpenStudy (nurali):

y=5x+4 The slope of this line is 5. The slope of a perpendicular line is the negative reciprical: -1/5 Now use the point-slope formula to find the equation: y - y1 = m (x - x1) y - 2 = -1/5 (x + 3) y-2 = -1/5x - 3/5 y = -1/5x-3/5+2 y = -1/5x+7/5

OpenStudy (anonymous):

so y = -1/5x+7/5 is the answer

OpenStudy (nurali):

Yes

OpenStudy (anonymous):

thank you

OpenStudy (nurali):

My Pleasure.